The gasoline tank of a car contains 50 liters of gasoline and alcohol, the alcohol comprising 25%. How much of the mixture must be drawn off and replaced by alcohol so that the tank contain a mixture of which 50% is alcohol?
Answer by Boreal(15235)

G+A=50 l
A=12.5 l
G=37.5 l
remove x liters and replace with x alcohol
each liter that is removed is 0.75 G and 0.25 A
37.5-0.75x=12.5-0.25x+x=12.5+0.75x
25=1.5x
x=16 2/3 l drawn off
after drawing off, there is 33 1/3 liters in tank and 1/4 is alcohol or 8.33 liters.
and 16 2/3 liters alcohol to the tank and have 50 liters again and 25 liters of alcohol or 50%
it is 16 2/3 liters.